OpenStudy (anonymous):

What is the probability that a randomly chosen point lies in the shaded region? (leave the answer as a fraction and leave pi in your answer)

OpenStudy (anonymous):

area of shaded region divided by total area

OpenStudy (anonymous):



OpenStudy (anonymous):

the triangle is shaded

OpenStudy (anonymous):

ok so the area of the circle is easy enough right? since the radius is 5 the area is \(\pi\times 5^2=25\pi\)

OpenStudy (anonymous):

Yes, I have that so far.

OpenStudy (anonymous):

and you have two triangles both with base 5 and height 5 so the total area of the two triangles is \(5\times 5=25\)

OpenStudy (anonymous):

take the ratio to get your answer

OpenStudy (anonymous):

So then isn't it just one..

OpenStudy (anonymous):

oh no the areas are not the same

OpenStudy (anonymous):

your shaded region has area 25, total circle has area \(25\pi\)

OpenStudy (anonymous):

oh wait never mind... so it's 25pi over 25?

OpenStudy (anonymous):

no other way around shaded area over total area

OpenStudy (anonymous):

so 25 over 25pi?

OpenStudy (anonymous):

don't forget probability is a number between 0 and 1 right cancel the 25

OpenStudy (anonymous):

1pi? hahha

OpenStudy (anonymous):

\(\frac{25}{25\pi}=\frac{1}{\pi}\)

OpenStudy (anonymous):

not \(1\pi\) but rather \(\frac{1}{\pi}\)

OpenStudy (anonymous):

Okay thank you!

OpenStudy (anonymous):

yw